# DMol: A Schedule-Driven Diffusion Model for Highly Efficient and Versatile Molecule Generation

### v1

@@ Abstract
We introduce a new graph diffusion model for small molecule generation, DMol, which outperforms the state-of-the-art DiGress model in terms of validity by roughly $1.5\%$ across all benchmarking datasets while reducing the number of diffusion steps by at least $10$-fold, and the running time to roughly one half. The performance improvements are a result of a careful change in the objective function and a “graph noise” scheduling approach which, at each diffusion step, allows one to only change a subset of nodes of varying size in the molecule graph. Another relevant property of the method is that it can be easily combined with junction-tree-like graph representations that arise by compressing a collection of relevant ring structures into supernodes. Unlike classical junction-tree techniques that involve VAEs and require complicated reconstruction steps, compressed DMol directly performs graph diffusion on a graph that compresses only a carefully selected set of frequent carbon rings into supernodes, which results in straightforward sample generation. This compressed DMol method offers additional validity improvements over generic DMol of roughly $2\%$, increases the novelty of the method, and further improves the running time due to reductions in the graph size.
Machine Learning, ICML
@@ 1 Introduction
Graphs are fundamental data structures used to model a wide range of complex interactions, including molecules and drugs, biological pathways, and social and co-purchase networks (Wu et al., 2021; Zhang et al., 2024; Yang et al., 2024; Waikhom & Patgiri, 2021). Therefore, the ability to generate graphs that accurately capture domain-specific distributions is of great significance (Tsai et al., 2023; You et al., 2018; Gamage et al., 2020). While generative models for images and texts are already being used with great success (Dhariwal & Nichol, 2021; Ho et al., 2022; Waikhom & Patgiri, 2021), graph generation remains a challenging frontier due to the discrete nature of the models, the need for permutation invariance and other distinctive properties (Velikonivtsev et al., 2024; Wang et al., 2023). This is particularly the case when trying to generate graph samples that mimic and innovate molecular compounds (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024; You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019; Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021b), in which case the generated samples can fail to obey necessary biochemical and physical constraints.
@@
One way to resolve some of these challenges is to use diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Fan et al., 2023; Chen et al., 2024; Trippe et al., 2023; Haefeli et al., 2023; Niu et al., 2020b; Huang et al., 2022; Liu et al., 2024c) which rely on a gradual noise injection process that perturbs the input samples and then reverses the process using a learned denoising network. In the context of small molecule drug generation, existing methods focus on discrete diffusion in order to avoid the loss of key structural properties that arise when embedding graph structures into continuous spaces (Niu et al., 2020a; Jo et al., 2022b). Discrete diffusion models incorporate “transition mechanisms” that explicitly consider the categorical attributes of nodes and edges, ensuring structural fine-graining and integrity during the noise injection process (Liu et al., 2024b; Xu et al., 2024). They also allow for continuous and discrete denoising steps that diversify the outputs (Huang et al., 2022; Vignac et al., 2023; Jo et al., 2022a). However, they also typically require a very large number of diffusion steps, resulting in large running times and extremely slow graph generation (Kong et al., 2023). More importantly, they lead to structures that have low chemical validity even when measured through highly coarse SMILES nomenclature constraints (Hoogeboom et al., 2022; Liu et al., 2021).
@@
The state-of-the-art graph diffusion method for generating small molecules, DiGress (Vignac et al., 2023), relies on a “cosine” noise injection mechanism that controls the amount of perturbations during the different diffusion steps. In the process, every node and edge in the graph is allowed to be modified at each step. DiGress evaluates the generated drug samples through validity, novelty, and uniqueness scores that provide a limited measure of the utility of molecules. Unfortunately, its running time is excessive and almost all samples generated in a run have unacceptably low biochemical scores seldom tested in the literature. Furthermore, in its current form, and unlike some non-diffusion based methods (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024; You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019), DiGress cannot preserve the structures and frequencies of important network motifs (e.g., carbon rings, which are the most important building blocks of molecules).
@@
We describe a new small molecule drug diffusion model, (Compressed) DMol, which significantly outperforms DiGress across all benchmarking datasets in terms of running time, SMILES validity, and novelty scores and at the same time, allows for network motif preservation via compressed representations of subsets of rings. Its main features are:
@@
1. DMol directly relates the diffusion time-steps with the number of nodes and edges that are allowed to undergo state changes in that step. This approach effectively increases the learning rate of the model as it learns increasingly larger submodels in a gradual manner. It also significantly reduces the time required for molecule generation. Furthermore, it uses a new objective function that penalizes mismatches in the count of nodes and edges during forward and backward steps. Unlike DiGress, DMol also exhibits a more flexible control of the noise schedule/distribution.
2. DMol can operate on graphs which “compress” important node motifs into supernodes. In compressed DMol, the key idea is to combine diffusion models with compressed motif representations of molecules akin to the ones used in JT-VAE (Jin et al., 2018). But unlike JT-VAE which requires sophisticated and time-consuming decoding, compressed DMol only uses a small number of carbon ring structures as super nodes and performs diffusion on the compacted graphs directly. Since the rings are chosen in such a way that only one type of bond is possible during reconstruction, the decoding is extremely simple. Ring compression differs from ring freezing as it still allows one ring structure to be changed to another. The validity improvement of compressed DMol over DiGress is $>3\%$ over all benchmarking datasets while the number of diffusion steps remains an order of magnitude smaller.
3. DMol also uniformly improves the performance of DiGress under previously unexplored validity metrics, such as ChEMBL likeness (which indicates how likely the generated molecule is to have similar biological activity as an already known drug in the ChEMBL molecule repository) and QED (quantitative estimate of drug-likeness), which is integrated into RdKit and measures chemical “desirability” (Landrum et al., 2006).
@@ 2 Related Works
Viewing small molecules as graphs has inspired a myriad of generative models for drug design, including JT-VAE (Jin et al., 2018), MoFlow (Zang & Wang, 2020), CDGS (Huang et al., 2023), EDM (Hoogeboom et al., 2022), etc. (Cornet et al., 2024; Shi et al., 2021; Xu et al., 2024; Chen et al., 2023; Wu et al., 2022; Bao et al., 2023). Furthermore, numerous surveys on the subject are readily available (Tang et al., 2024; Du et al., 2022; Yang et al., 2024). We partition and review prior works on generative models for drug discovery based on the methodology used.
2.1 Traditional Methods
@@
Traditional approaches to graph-based molecular generation relied on the use of Variational Autoencoders (VAEs) as generative models (Jin et al., 2018; Simonovsky & Komodakis, 2018; Mitton et al., 2021; Vignac & Frossard, 2021; Gómez-Bombarelli et al., 2018; Kwon et al., 2020; Wu et al., 2024). VAEs comprise an encoder that maps the input graph into a latent space and a decoder that reconstructs the graph from the latent embedding. Many methods also involve autoregressive models, which generate molecular graphs by sequentially predicting the next element in the sequence based on previous outputs (You et al., 2018; Liao et al., 2019; Segler et al., 2018; Popova et al., 2019). On the other hand, flow-based methods leverage the concept of normalizing flows for graph generation (Madhawa et al., 2019; Zang & Wang, 2020; Luo et al., 2021b; Lippe & Gavves, 2021; Shi* et al., 2020). These techniques apply a series of invertible transformations to model complicated distributions based on simpler ones (such as a Gaussian). Additionally, Generative Adversarial Networks (GANs) approaches have also been used for molecule generation (De Cao & Kipf, 2018; Łukasz Maziarka et al., 2019; Martinkus et al., 2022a). GANs include a generator, which learns to create realistic molecular graphs, and a discriminator, which learns to differentiate between real and generated samples.
2.2 Diffusion-Based Methods
@@
Diffusion models for graph generation can be broadly categorized as follows. Denoising Diffusion Probabilistic Models (DDPM) utilize a Markov chain for the diffusion process (Vignac et al., 2023; Xu et al., 2022; Liu et al., 2024b; Xu et al., 2024; Liu et al., 2024a; Hoogeboom et al., 2022; Jo et al., 2024). Score-Based Generative Models (SGM) leverage score matching techniques to model the data distribution, and generate graphs by reversing a diffusion process guided by the score of the data distribution (Luo et al., 2021a; Chen et al., 2023; Wu et al., 2022). Furthermore, by replacing discrete time steps with continuous time, one can use stochastic differential equations to model the diffusion and denoising processes (Jo et al., 2022a; Lee et al., 2023; Huang et al., 2023; Bao et al., 2023). Furthermore, studies have shown that different noise schedules may significantly affect molecule generation quality (Shi et al., 2025; Nichol & Dhariwal, 2021). Comprehensive surveys on the subjects include (Mengchun Zhang et al., 2023; Fan et al., 2023; Chen et al., 2024).
@@
DiGress (Vignac et al., 2023) uses a denoising diffusion probabilistic model within a discrete state space and is considered one of the state-of-the-art models for molecular generation tasks. It represents graphs in the discrete space of node and edge attributes by assigning categorical labels to each node and edge. DiGress models the noise addition process as a Markov process, where each node and edge label evolves independently of the others. This assumption mirrors the approach in image-based diffusion models. Denoising is performed using a graph transformer network, trained by minimizing the cross-entropy loss between the predicted probabilities for nodes and edges and the ground-truth graph. The transformer network reconstructs a denoised graph from a noisy input. To generate new graphs, a noisy graph is first sampled according to the model’s limiting distribution, and the trained denoising transformer is then used to produce a graph with the desired properties. The DiGress model is discussed further in Section 5.
@@ 3 DMol Diffusion
3.1 Review of the Graph Diffusion Model
@@
More precisely, the encoding of a graph involves categorical information about the nodes $V$ and vertices $E$ of the graph $\mathcal{G}=(V,E).$ The assumptions are that each node belongs to one of $a$ classes and is represented as a one-hot vector in $\mathbb{R}^{a}$. The same is true for edges, which are assumed to belong to one of $b$ classes (with one class indicating the absence of the edge). We follow the notation from (Vignac et al., 2023), where $\textbf{x}_{i}$ is used to denote the one-hot encoding of the class of node $i$. The one-hot encodings are arranged in a node matrix $\bf{X}$ of dimensions $n\times a$. Again, in a similar manner, edges are encoded into a tensor $\bf{E}$ of dimension $n\times n\times b$. The Markovian state space for each node and edge is the set of classes of the nodes and edges, respectively. For simplicity, and with a slight abuse of notation, $\textbf{x}_{l}$ is used to denote the class (state) of node $l$, while $\textbf{e}_{ls}$ is used to denote the class (state) of edge $ls$.
@@
Diffusion is performed separately on nodes and edges, while the noisy perturbations in the forward process are governed by a state transition matrix. For node-level noise we use transition matrices $(\textbf{Q}_{X}^{1},\dots,\textbf{Q}_{X}^{T})$, where each entry $[\textbf{Q}_{X}^{t}]_{ij}$ defines the probability of an arbitrary node transitioning from state $i$ to state $j$ at time step $t$. A similar model is used for edges, where the transition matrices $(\textbf{Q}_{E}^{1},\dots,\textbf{Q}_{E}^{T})$ describe the probabilities $[\textbf{Q}_{E}^{t}]_{ij}$ of an arbitrary edge transitioning from state $i$ to state $j$. Formally, $[\textbf{Q}_{X}^{t}]_{ij}=q(x^{t}=j\mid x^{t-1}=i)$ and $[\textbf{Q}_{E}^{t}]_{ij}=q(e^{t}=j\mid e^{t-1}=i)$, where $x$ and $e$ correspond to a generic node and edge, while the superscript indicates the time stamp.
@@
Consequently, the transition probability at time $t$ is given by $q(\textbf{z}^{t}\mid\textbf{z}^{t-1})={\textbf{z}^{t-1}}\textbf{Q}^{t}$, where z stands for either x or e and the subscript of Q is set accordingly. Adding noise to a graph $\mathcal{G}^{t-1}$ at time $t-1$ results in a new graph $\mathcal{G}^{t}=(\textbf{X}^{t},\textbf{E}^{t}),$ and involves sampling each node and edge class from a categorical distribution succinctly described as $q(\mathcal{G}^{t}\mid\mathcal{G}^{t-1})=\big{(}\textbf{X}^{t-1}\textbf{Q}_{X}^{t},\textbf{E}^{t-1}\textbf{Q}_{E}^{t}\big{)}$ and $q(\mathcal{G}^{t}\mid\mathcal{G}^{0})=\big{(}\textbf{X}^{0}\overline{\textbf{Q}}_{X}^{t},\textbf{E}^{0}\overline{\textbf{Q}}_{E}^{t}\big{)}$, where $\overline{\textbf{Q}}_{X}^{t}=\textbf{Q}_{X}^{1}\dots\textbf{Q}_{X}^{t}$ and $\overline{\textbf{Q}}_{E}^{t}=\textbf{Q}_{E}^{1}\dots\textbf{Q}_{E}^{t}$. The state transition matrices are invertible, which ensures the reversibility of the noise model: the graphs at time $t$ and time 00 can be transformed into each other using the state transition matrix. Reversibility guarantees consistency between the forward and backward processes and simplifies denoising in the backward process.
@@
A common issue with most known diffusion models is the large number of steps needed for convergence (Hang et al., 2023; Wang et al., 2024). Although there are typically no strict time constraints for molecule generation, for large training sets the computational complexity may be overwhelming (see also the Results section). Hence, one important question is how to reduce the number of diffusion steps for graph generation, and at the same time, improve the utility of the model. These issues are addressed via our new DMol diffusion model.

### v3

# DMol: A Highly Efficient and Chemical Motif-Preserving Molecule Generation Platform

@@ Abstract
We introduce a new graph diffusion model for small drug molecule generation which simultaneously offers a $10$-fold reduction in the number of diffusion steps when compared to existing methods, preservation of small molecule graph motifs via motif compression, and an average $3\%$ improvement in SMILES validity over the DiGress model across all real-world molecule benchmarking datasets. Furthermore, our approach outperforms the state-of-the-art DeFoG method with respect to motif-conservation by roughly $4\%$, as evidenced by high ChEMBL-likeness, QED and newly introduced shingles distance scores. The key ideas behind the approach are to use a combination of deterministic and random subgraph perturbations, so that the node and edge noise schedules are codependent; to modify the loss function of the training process in order to exploit the deterministic component of the schedule; and, to “compress” a collection of highly relevant carbon ring and other motif structures into supernodes in a way that allows for simple subsequent integration into the molecular scaffold [^1] The link to the code can be found at: https://github.com/liekon/Discrete-Graph-Generation.
@@ 1 Introduction
Graphs are fundamental data structures used to model a wide range of complex interactions, including molecules and drugs, biological pathways, and social and co-purchase networks Wu et al. (2021); Zhang et al. (2024); Yang et al. (2024); Waikhom and Patgiri (2021). Hence, the ability to generate graphs that accurately capture domain-specific distributions is of great significance Tsai et al. (2023); You et al. (2018); Gamage et al. (2020). While generative models for images and texts are being used with great success Dhariwal and Nichol (2021); Ho et al. (2022); Waikhom and Patgiri (2021), graph generation remains a challenging frontier due to the discrete nature of the models, the need for permutation invariance and other distinctive properties Velikonivtsev et al. (2024); Wang et al. (2023). This is particularly the case when trying to generate graph samples that innovate molecular compounds Jin et al. (2018); Mitton et al. (2021); Vignac and Frossard (2021); Gómez-Bombarelli et al. (2018); Kwon et al. (2020); Wu et al. (2024); You et al. (2018); Liao et al. (2019); Segler et al. (2018); Popova et al. (2019); Madhawa et al. (2019); Zang and Wang (2020), in which case the generated samples can fail to obey necessary biochemical and physical constraints and preserve important motif compounds.
@@
One way to address these challenges is to use diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Fan et al. (2023); Chen et al. (2024); Trippe et al. (2023); Haefeli et al. (2023); Niu et al. (2020b); Huang et al. (2022); Liu et al. (2024c) which rely on a gradual noise injection process that perturbs the input samples and then reverses the process using a learned denoising network. In the context of small molecule drug generation, existing methods focus on discrete diffusion in order to avoid the loss of key structural properties that arise when embedding graph structures into continuous spaces Niu et al. (2020a); Jo et al. (2022b). Discrete diffusion models incorporate “transition mechanisms” that explicitly consider the categorical attributes of nodes and edges, ensuring fine-grained structural changes during the noise injection process Liu et al. (2024b); Xu et al. (2024a). They also allow for continuous and discrete denoising steps that diversify the outputs Huang et al. (2022); Vignac et al. (2023); Jo et al. (2022a). However, they also typically require a very large number of diffusion steps, resulting in large running times and extremely slow graph generation Kong et al. (2023). More importantly, they lead to structures that have low chemical validity even when measured through exceptionally coarse SMILES nomenclature constraints Hoogeboom et al. (2022); Liu et al. (2021). Most learning methods consequently fail to produce molecules that can be chemically synthesized or that can properly dock on target proteins.
@@
One well-known graph diffusion method, DiGress Vignac et al. (2023), relies on a cosine noise injection mechanism that controls the amount of perturbations during the different diffusion steps. In the process, every node and edge in the graph is allowed to be modified at each step. DiGress evaluates the generated drug sample quality through SMILES validity, novelty, and uniqueness all of which provide a highly limited measure of the utility of molecules. Also, the training and sampling time of the method is excessive. Furthermore, in its current form, and unlike some nondiffusion based methods Jin et al. (2018); Simonovsky and Komodakis (2018); Mitton et al. (2021); Vignac and Frossard (2021); Gómez-Bombarelli et al. (2018); Kwon et al. (2020); Wu et al. (2024); You et al. (2018); Liao et al. (2019); Segler et al. (2018); Popova et al. (2019), DiGress cannot preserve the structures and frequencies of important drug network motifs (including carbon rings, which are the most important building blocks of organic molecules).
@@
We describe a new small molecule drug diffusion model, Diffusion Models for Molecular Motifs, DMol, which significantly outperforms DiGress across all benchmarking datasets in terms of running time, SMILES validity, and novelty scores and at the same time, allows for network motif preservation via compressed representations of subsets of motifs. At the same time, the DMol improves the more informative chemical validity scores when compared to the very recent state-of-the-art DeFoG Qin et al. (2025) method by roughly $4\%$, since the latter cannot guarantee preservation of motifs. Its main features are as follows:
@@
1. DMol directly relates the diffusion time-steps with the number of nodes and edges that are allowed to undergo state changes in that step. This approach effectively increases the learning rate of the model as it learns increasingly larger submodels in a gradual manner. It also significantly reduces the time required for molecule generation, leading to an order of magnitude reduction in the number of steps required by diffusion models.
2. The number of nodes and edges that DMol perturbs at each step is deterministic, but the nodes themselves are selected at random. Furthermore, the perturbed edges are confined to the subgraph induced by the selected vertices, which couples the node and edge perturbations and leads to better preservation of subgraph structures.
3. DMol uses new penalty terms in the objective function that penalize mismatches in the deterministic counts of nodes and edges perturbed during the forward and backward steps. Such penalties lead to poor results when the number of perturbed nodes is random, as in DiGress and other methods, and motivates the use of a mixed deterministic/random noise schedule. This allows for further “forced motif” preservation and a more flexible control of the noise schedule/distribution, as evidenced by a straightforward theoretical analysis.
4. One of the most important features of DMol is motif compression which allows one to convert chemically important node motifs into supernodes that are diffused similarly to regular atom nodes. The chosen motifs (which are mostly carbon rings) have the property that they can always be reattached to the molecular scaffold, and the number of such molecules is bounded to retain computational savings guaranteed by the first three innovations. Due to their careful chemical selection, only one type of bond is possible during reconstruction, and “decoding” is extremely simple. Motif compression differs both from JT-VAEs Jin et al. (2019) and ring freezing proposed in DiGress (which failed to perform in a satisfactory manner), since in our case one is still allowed to directly replace one valid motif structure by a single atom or another valid motif structure (in addition, the choices of substructures used differ widely).
As a result, DMol improves the SMILES validity of DiGress by roughly $3\%$ over all real-world molecular benchmarking datasets while keeping the number of diffusion steps at an order that is one magnitude smaller. At the same time, the significantly more meaningful chemical likeness scores (e.g., RDKit’s Landrum et al. (2006) ChEMBL likeness, QED and shingle distances, which indicate how likely the generated molecule is to have desired chemical motifs and similar biological activity properties) of DMol are consistently better than those of DiGress. These scores for DMol are roughly $4\%$ higher, on average, than those of the recent state-of-the-art flow matching method DeFoG Qin et al. (2025), since the latter does not tend to preserve chemical motifs.
As a final remark, it is important to point out that in the biochemistry literature, it has been long recognized that SMILES validity optimization may not be as practically relevant from the chemical synthesis/properties point of few; motif, and especially carbon ring structure preservation, is deemed much more important, but tangible evaluation metrics for true chemical validity are still lacking. See for example the recent work Skinnider (2024) that discusses pros and cons of SMILES validity.
@@ 2 Related Works
Viewing small molecules as graphs has inspired a myriad of generative models for drug design, including JT-VAE Jin et al. (2018), MoFlow Zang and Wang (2020), CDGS Huang et al. (2023), EDM Hoogeboom et al. (2022), etc Cornet et al. (2024); Shi et al. (2021); Xu et al. (2024a); Chen et al. (2023); Wu et al. (2022); Bao et al. (2023); You et al. (2018); Liao et al. (2019); Mercado et al. (2021); Segler et al. (2018); Kwon et al. (2020); Jensen (2019); Jo et al. (2024).
Furthermore, numerous surveys on the subject are readily available Tang et al. (2024); Du et al. (2022); Yang et al. (2024). We partition and review prior works on generative models for drug discovery based on the methodology used.
@@
Traditional Methods. Traditional approaches to graph-based molecular generation relied on the use of Variational Autoencoders (VAEs) as generative models Jin et al. (2018); Simonovsky and Komodakis (2018); Mitton et al. (2021); Vignac and Frossard (2021); Gómez-Bombarelli et al. (2018); Kwon et al. (2020); Wu et al. (2024). VAEs comprise an encoder that maps the input graph into a latent space and a decoder that reconstructs the graph from the latent embedding. Many methods also involve autoregressive models, which generate molecular graphs by sequentially predicting the next element in the sequence based on previous outputs You et al. (2018); Liao et al. (2019); Segler et al. (2018); Popova et al. (2019). On the other hand, flow-based methods leverage the concept of normalizing flows for graph generation Madhawa et al. (2019); Zang and Wang (2020); Luo et al. (2021b); Lippe and Gavves (2021); Shi* et al. (2020). These techniques apply a series of invertible transformations to model complicated distributions based on simpler ones (such as a Gaussian). Additionally, Generative Adversarial Networks (GANs) approaches have also been used for molecule generation De Cao and Kipf (2018); Łukasz Maziarka et al. (2019); Martinkus et al. (2022a). GANs include a generator, which learns to create realistic molecular graphs, and a discriminator, which learns to differentiate between real and generated samples.
@@
Diffusion-Based Methods. Diffusion models for graph generation can be broadly categorized as follows. Denoising Diffusion Probabilistic Models (DDPM) utilize a Markov chain for the diffusion process Vignac et al. (2023); Xu et al. (2022); Liu et al. (2024b); Xu et al. (2024a); Liu et al. (2024a); Hoogeboom et al. (2022); Jo et al. (2024). Score-Based Generative Models (SGM) leverage score matching techniques to model the data distribution, and generate graphs by reversing a diffusion process guided by the score of the data distribution Luo et al. (2021a); Chen et al. (2023); Wu et al. (2022). Furthermore, by replacing discrete time steps with continuous time, one can use stochastic differential equations to model the diffusion and denoising processes Jo et al. (2022a); Lee et al. (2023); Huang et al. (2023); Bao et al. (2023). Furthermore, studies have shown that different noise schedules may significantly affect molecule generation quality Shi et al. (2025); Nichol and Dhariwal (2021). Comprehensive surveys on the subjects include Mengchun Zhang et al. (2023); Fan et al. (2023); Chen et al. (2024).
@@
DiGress Vignac et al. (2023) uses a denoising diffusion probabilistic model within a discrete state space and is considered one of the practically effective models for molecular generation tasks. It represents graphs in the discrete space of node and edge attributes by assigning categorical labels to each node and edge. DiGress models the noise addition process as a Markov process, where each node and edge label evolves independently of the others. This assumption mirrors the approach in image-based diffusion models. Denoising is performed using a graph transformer network, trained by minimizing the cross-entropy loss between the predicted probabilities for nodes and edges and the ground-truth graph. The transformer network reconstructs a denoised graph from a noisy input. To generate new graphs, a noisy graph is first sampled according to the model’s limiting distribution, and the trained denoising transformer is then used to produce a graph with the desired properties. The DiGress model is discussed further in Section B.
DisCo Xu et al. (2024b) uses discrete-state continuous-time diffusion, addressing limitations of discrete-time approaches like DiGress. The model formulates graph generation as a continuous-time Markov Chain (CTMC) that preserves the discrete nature of graph-structured data while enabling flexible sampling trade-offs. Unlike discrete-time models with fixed sampling steps, DisCo can adjust sampling steps after training without retraining the model. DisCo enables numerical approaches like $\tau$-leaping for efficient simulation of the reverse process, allowing practitioners to dynamically balance generation quality and computational efficiency during inference. Despite its innovations, DisCo suffers from quadratic computational complexity with respect to the number of nodes, limiting its scalability for generating large graphs. Moreover, DisCo reduces the number of sampling steps at the cost of degraded generation quality.
Flow Matching-based Methods. Flow Matching (FM) has recently emerged as an alternative to diffusion models for generative tasks Lipman et al. (2023); Dao et al. (2023). FM frameworks define continuous probability paths between a simple prior distribution and the target data distribution, offering comparable performance and efficiency compared to diffusion-based approaches. To address discrete state spaces, Discrete Flow Matching (DFM) formulations have been developed Gat et al. (2024), providing a streamlined approach with more flexible sampling procedures. DeFoG Qin et al. (2025) extends the DFM framework specifically for graph generation tasks. It employs a linear interpolation noising process and a CTMC-based denoising process while preserving the inherent symmetries of graphs. The key innovation in DeFoG is its disentanglement of training and sampling stages, which enables independent optimization of each component. Despite its innovation, DeFoG still requires dataset-specific hyperparameter tuning for optimal sampling strategies, making its adoption potentially challenging for new graph domains without extensive experimentation. Importantly, experiments to be described in this work reveal that DeFoG under-performs on key biochemical motif molecular property metrics such as QED.
@@ 3 DMol Diffusion
3.1 Review of Standard Graph Diffusion Models
@@
More precisely, the encoding of a graph involves categorical information about the nodes $V$ and vertices $E$ of the graph $\mathcal{G}=(V,E).$ The assumptions are that each node belongs to one of $a$ classes and is represented as a one-hot vector in $\mathbb{R}^{a}$. The same is true for edges, which are assumed to belong to one of $b$ classes (with one class indicating the absence of the edge). We follow the notation from Vignac et al. (2023), where $\textbf{x}_{i}$ is used to denote the one-hot encoding of the class of node $i$. The one-hot encodings are arranged in a node matrix $\bf{X}$ of dimensions $n\times a$. Again, in a similar manner, edges are encoded into a tensor $\bf{E}$ of dimension $n\times n\times b$. The Markovian state space for each node and edge is the set of classes of the nodes and edges, respectively. With a slight abuse of notation, $\textbf{x}_{l}$ stands for the class (state) of node $l$, while $\textbf{e}_{ls}$ stands for the class (state) of edge $ls$.
@@
Diffusion is performed separately on nodes and edges, while the noisy perturbations in the forward process are governed by a state transition matrix. For node-level noise we use transition matrices $(\textbf{Q}_{X}^{1},\dots,\textbf{Q}_{X}^{T})$, where each entry $\textbf{Q}_{X}^{t}]_{ij}$ defines the probability of an arbitrary node transitioning from state $i$ to state $j$ at time step $t$. A similar model is used for edges, where the transition matrices $(\textbf{Q}_{E}^{1},\dots,\textbf{Q}_{E}^{T})$ describe the probabilities $\textbf{Q}_{E}^{t}]_{ij}$ of an arbitrary edge transitioning from state $i$ to state $j$. Formally, $\textbf{Q}_{X}^{t}]_{ij}=q(x^{t}=j\mid x^{t-1}=i)$ and $\textbf{Q}_{E}^{t}]_{ij}=q(e^{t}=j\mid e^{t-1}=i)$, where $x$ and $e$ correspond to a generic node and edge, while the superscript indicates the time stamp.
@@
Consequently, the transition probability at time $t$ is given by $q(\textbf{z}^{t}\mid\textbf{z}^{t-1})={\textbf{z}^{t-1}}\textbf{Q}^{t}$, where z stands for either x or e and the subscript of Q is set accordingly. Adding noise to a graph $\mathcal{G}^{t-1}$ at time t1t-1 results in a new graph $\mathcal{G}^{t}=(\textbf{X}^{t},\textbf{E}^{t}),$ and involves sampling each node and edge class from a categorical distribution succinctly described as $q(\mathcal{G}^{t}\mid\mathcal{G}^{t-1})=\big(\textbf{X}^{t-1}\textbf{Q}_{X}^{t},\textbf{E}^{t-1}\textbf{Q}_{E}^{t}\big)$ and $q(\mathcal{G}^{t}\mid\mathcal{G}^{0})=\big(\textbf{X}^{0}\overline{\textbf{Q}}_{X}^{t},\textbf{E}^{0}\overline{\textbf{Q}}_{E}^{t}\big)$, where $\overline{\textbf{Q}}_{X}^{t}=\textbf{Q}_{X}^{1}\dots\textbf{Q}_{X}^{t}$ and $\overline{\textbf{Q}}_{E}^{t}=\textbf{Q}_{E}^{1}\dots\textbf{Q}_{E}^{t}$. The state transition matrices are invertible, which ensures the reversibility of the noise model: the graphs at time $t$ and time 0 can be transformed into each other using the state transition matrix. Reversibility guarantees consistency between the forward and backward processes and simplifies backward denoising.
@@
A common issue with most known diffusion models is the large number of steps needed for convergence Hang et al. (2023); Wang et al. (2024). For large training sets the computational complexity may be overwhelming (see our Results section). Hence, the first important question is how to reduce the number of diffusion steps for graph generation, and at the same time, preserve the chemical utility of the model. These issues are addressed by DMol.